You have a lot you believe to be 157' x 157'. After lot is surveyed you discover that one side is 2' shorter and the other side is 2' longer. How much less area do you have than originally believed?
Answer by stanbon(75887)

(157-2)(157+2) = 157^2-2^2 = (area you believed)-4
------------
You have 4 square feet less area
==============
Cheers,
Stan H.